OpenStudy (anonymous):

4(7x - 1) = 3(2x + 1) - 4(x + 6)

OpenStudy (tennistar):

first distribute and you get 21x -4 = 6x + 3 - 4x -24

OpenStudy (tennistar):

now subtract -4 from both sides

OpenStudy (tennistar):

I mean add oops

OpenStudy (tennistar):

:)

OpenStudy (tennistar):

21x = 6x + 7 - 4x - 24 and subtract 6x from both sides

OpenStudy (tennistar):

so you get 15x = 7 -4x-24 and add 4x and get 19x = 7-24

OpenStudy (anonymous):

7*4 is 28

OpenStudy (tennistar):

yeah but its a minus

OpenStudy (tennistar):

19x = 17 if I solved correct

OpenStudy (tennistar):

x = 17/19 or seventeen over nineteen

OpenStudy (anonymous):

a minus doesnt make 4*7 =21

OpenStudy (anonymous):

I want to say it would be this: 28x - 4 = 6x + 3 - 4x - 24 -- Multiplied 26x = -17 -- Simplified x = -17/26

OpenStudy (anonymous):

thats what i got the third time thanks much!